OpenStudy (anonymous):

HELP!! Has the polynomial below been factored correctly and completely? x5 - 16x3 + 8x2 - 128 x3(x2 - 16) + 8(x2 - 16) (x2 -16)(x3 + 8)

OpenStudy (tkhunny):

Or not. Keep trying.

jhonyy9 (jhonyy9):

correctly and completely so completely not is sure because these factors can being again factorized ,like a^2 -b^2 = (a-b)(a+b) and a^3 +b^3 = (a+b)(a^2 -ab +b^2) hope these will help you ending it easy and right